Solution Dilution Learning Goal To understand how to prepare a diute solution and determine the concentration or volume of the initial or Initial Solution O Type here to search Part B what is the sna volume in militers when o644 Lof a 31.5% (m/) solution is duted to 248% (m/v)? Express your answer with the appropriate units. v Value Units Hits My Answers Give Up Review Part Part C A 510 mL NaCI solution is dauted to a volume of 1,34 Land a concontration of 6.00 M. What was tho initial conoontation? Express your answer with the appropriate units, M Value Units Hints My Answers Give up Review Part
Answer #1

Part B

Use the formula C1V1 =C2V2

Where, C1 = initial concentration = 31.5 %

V1 = initial volume = 0.644 L = 644 ml

C2 = final concentration = 24.8 %

V2 = final volume = ?

C1V1 =C2V2 Thus,

V2 = C1V1 /C2

Substitute value in above equation

V2 = 31.5 \times 644 / 24.8 = 818 ml

final volume = 818 ml

Part C

Use the formula C1V1 = C2V2

Where C1 = initial concentration = ?

V1 = initial volume = 510 ml

C2 = final concentration = 6.00M

V2 = final volume = 1.34 L = 1340 ml

Substitute the value in formula

C1V1 = C2V2

C1 = C2V2/V1

= 6.00 \times 1340 / 510 = 15.76

C1 = 15.76 M

initial concentration of NaCl = 15.76 M